OpenStudy (anonymous):

A skier takes off from the top of the ski run and reaches a speed of 24m/s in 4 seconds. What is the skier's acceleration? (Assume that starting velocity is = 0) Given: Equation & Work: vi= vf= t= a= answer ____________

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

@iGreen ?????? Anything?

OpenStudy (anonymous):

@TheSmartOne please?

OpenStudy (anonymous):

@SolomonZelman please help

OpenStudy (anonymous):

\[\frac{ dv }{ dt } =\frac{ 24 }{ 4 }\]

OpenStudy (anonymous):

so they gain a speed of 6m/s every second, which is 24m/s in 4 seconds

OpenStudy (anonymous):

as we know that v=d/t velocity =distance /time so distance =velocity*time d=24 *4=>96m and frm above eq its simple a =6 metre per second square